# Bokeh Diffusion: Defocus Blur Control in Text-to-Image Diffusion Models

### v1

@@ 2 Related Work
Most existing works on camera control in diffusion models focus on extrinsic parameters, such as camera position and viewing angle [18, 17, 9], or trajectory in video generation [44, 13, 38, 46, 47]. In contrast, intrinsic camera control has received comparatively little attention. Curved Diffusion [39] explores conditioning T2I diffusion on diverse optical geometries to simulate lens-induced distortions, such as fisheye and panoramic effects. Additionally, some methods achieve zoom effects by implicitly modeling focal length [43, 9].
Generative Photography [50] is a concurrent work that disentangles intrinsic camera control by modeling different camera parameters across the temporal axis of a text-to-video diffusion model. In contrast, our method remains within a T2I paradigm and learns to transfer the depth-of-field from one image while anchoring the scene structure of another. This allows our approach to be directly applied to widely used T2I models and avoids the overhead of constantly generating several frames at inference time.
@@ 3.2 Leveraging In-the-wild Image Augmentation
Even so, the scarcity of truly paired images with multiple bokeh states per scene remains a critical bottleneck. To overcome this, we synthesize additional defocus examples by rendering artificial blur on nearly all-in-focus images. Specifically, we estimate a per-pixel disparity map and then sample a blur parameter $K$ to generate contrastive “focus–defocus” states of the same image. Such synthetic data helps the model avoid local minima: it sees direct evidence that the same object layout can be rendered sharply or with shallow focus, thus faster learning to separate bokeh from scene semantics.
@@ 4.2 Scene-consistent Defocus Blur Control
  • Accuracy: We measure how each method’s bokeh aligns with the target blur parameter. Specifically, we compute the average Laplacian variance trend for images in a scene under different blur conditions, then compute Pearson correlation with the reference from BokehMe (LVCorr). A higher correlation indicates more faithful defocus control. Additionally, we employ GPT-4o [27] to estimate the bokeh level of a batch of images and calculate the Pearson correlation with its conditioning sequence.
@@
Method Accuracy Consistency
@@
Table 1 shows that our approach outperforms all baselines by a significant margin in both accuracy and consistency. Standard T2I diffusion models struggle to achieve reliable control over depth-of-field, as reflected in their low correlation scores for blur trends. Furthermore, modifying the condition frequently alters the scene, as evidenced by the consistency scores. Figure 6 offers a qualitative comparison: while the vanilla T2I models often alter the subject arrangement or fail to match the intended blur strength, our method faithfully changes bokeh levels while preserving the scene. Finally, image–text alignment measured via CLIP embeddings yields comparable scores (around 0.330) for our method and the base SD1.5 model, indicating that our additional modules do not compromise the generative quality of the underlying model.
@@
Our framework accommodates two usage scenarios: (i) Unbound bokeh generation, where a user can directly sample an image at the desired blur level (see supplementary Fig. 12 for examples). (ii) Grounded bokeh generation, where a pivot image is chosen to anchor the scene whose bokeh the user wants to adjust. Since we randomly pick a pivot in each contrastive batch during training, the model learns to either add or remove blur relative to that pivot. Consequently, at inference time we can choose any pivot image with a given bokeh level and request a new target level. Figure 7 demonstrates how using either a sharper or blurrier pivot yields results matching the requested bokeh parameter without changing contents.
@@ Appendix C Additional Visual Results
Figure 12: Visual results for unbound bokeh-conditioned generation.
Figure 11 shows further scene-consistent sharp/bokeh generated pairs from our model. In Figure 12, we present visual results for unbound bokeh-conditioned generation.

### v3

# BoookkkkkeeeeehhhhhDiffusion: Defocus Blur Control in Text-to-Image Diffusion Models

@@ 2 Related Work
Most existing works on camera control in diffusion models focus on extrinsic parameters, such as camera position and viewing angle [19, 18, 9], or trajectory in video generation [47, 14, 41, 49, 50]. In contrast, intrinsic camera control has received comparatively little attention. Curved Diffusion [42] explores conditioning T2I diffusion on diverse optical geometries to simulate lens-induced distortions, such as fisheye and panoramic effects. Additionally, some methods achieve zoom effects by implicitly modeling focal length [46, 9], while other approaches model some explicit camera effects without ensuring scene consistency [11, 28].
Generative Photography [53] is a recent work that disentangles intrinsic camera control along the temporal axis of a text-to-video model. In contrast, our method remains within a T2I paradigm and learns to transfer the depth-of-field from one image while anchoring the scene structure of another. This allows our approach to be directly applied to widely used T2I models and avoids the overhead of constantly generating several frames at inference time.
@@ 3.2 Leveraging In-the-wild Image Augmentation
Even so, the scarcity of truly paired images with multiple bokeh states per scene remains a critical bottleneck. To overcome this, we synthesize additional defocus examples by rendering artificial blur on nearly all-in-focus images [53]. Specifically, we estimate a per-pixel disparity map and then sample a blur parameter $K$ to generate contrastive “focus–defocus” states of the same image. Such synthetic data helps the model avoid local minima: it sees direct evidence that the same object layout can be rendered sharply or with shallow focus, thus faster learning to separate bokeh from scene semantics.
@@ 4.2 Scene-consistent Defocus Blur Control
  • Accuracy: We measure how each method’s bokeh aligns with the target blur parameter via CorrCoef [53]. Specifically, we compute the average Laplacian variance trend for images in a scene under different blur conditions, then compute Pearson correlation with the reference (LVCorr). A higher correlation indicates more faithful defocus control. Additionally, we employ GPT-4o [29] to estimate the bokeh level of a batch of images and calculate the Pearson correlation with its conditioning sequence.
@@
Method Accuracy (CorrCoef) Consistency
@@
Table 1 shows that our approach outperforms all baselines by a significant margin in both accuracy and consistency. Standard T2I diffusion models struggle to achieve reliable control over depth-of-field, as reflected in their low correlation scores for blur trends. Furthermore, modifying the condition frequently alters the scene, as evidenced by the consistency scores. Figure 6 offers a qualitative comparison: while the vanilla T2I models often alter the subject arrangement or fail to match the intended blur strength, our method faithfully changes bokeh levels while preserving the scene. Finally, image–text alignment measured via CLIP [34] embeddings yields comparable scores (around 0.330) for our method and the base SD1.5 model, indicating that our additional modules do not compromise the generative quality of the underlying model.
@@
Our framework accommodates two usage scenarios: (i) Unbounded bokeh generation, where a user can directly sample an image at the desired blur level (see supplementary Fig. 12). (ii) Grounded bokeh generation, where a pivot image is chosen to anchor the scene whose bokeh the user wants to adjust. Since we randomly pick a pivot in each contrastive batch during training, the model learns to either add or remove blur relative to that pivot. Consequently, at inference time we can choose any pivot image with a given bokeh level and request a new target level. Figure 7 demonstrates how using either a sharper or blurrier pivot yields results matching the requested bokeh parameter without changing contents.
@@ Appendix C Additional Visual Results
Figure 12: Visual results for unbounded bokeh-conditioned generation.
Figure 11 shows further scene-consistent sharp/bokeh generated pairs from our model. In Figure 12, we present visual results for unbounded bokeh-conditioned generation.